OpenStudy (anonymous):

Which is the slope of the line with equation 4x 2y = 4? Option A: 2 Option B: Option C: Option D: 2

OpenStudy (anonymous):

Is there a minus sign or plus sign between the 4x and 2y terms? If a minus sign, then: 4x - 2y = 4 -2y = -4x + 4 y = 2x - 2 So, the slope would be 2, or 2 to 1, or 2/1. If a plus sign, then: 4x + 2y = 4 2y = -4x + 4 y = -2x +2 So, the slope would be -2, or -2 to 1, or -2/1.

OpenStudy (anonymous):

I got it! Thank you (:

OpenStudy (anonymous):

You're welcome!! Xo!